Circular No. 7896

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 BV = 2001 CB_{40} (YEUNG) An apparently asteroidal object discovered by William Kwong Yeung,

Benson, AZ, on CCD images taken with a 0.45-m reflector near Apache Peak on Jan. 21.5 UT (discovery observation below taken from MPS 47535) was identified by the Minor Planet Center with additional apparently asteroidal observations (including some in 1998 and 2000–2001) made at several observatories through its routine processing. Noting the unusual nature of its orbit, T. Spahr obtained unfiltered CCD observations with M. Calkins at the 1.2-m reflector on Mount Hopkins on May 5, 6, and 7 that show the object (m_{1} ∼ 17) larger than nearby stars of similar brightness and with a persistent faint tail ∼ 5′′ long in p.a. 315^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 21.49201  15^{h}03^{m}44.^{s}64  −13^{o}47′01′′.1  20.4

Astrometry and elliptical orbital elements appear on MPEC 2002-J34:

Epoch = 2002 Mar. 27.0 TT

T = 2002 Mar. 11.05812 TT  ω = 178.^{o}77060
e = 0.3611535  Ω = 40.14116 }2000.0
q = 2.2444617 AU  i = 11.51510

a = 3.5133036 AU n^{o} = 0.14966855 P = 6.585 years

IM NORMAE C. E. Woodward, University of Minnesota; B. Miller, Gemini Observa-

tory; and S. G. Starrfield, Arizona State University, report that mediumresolution blue spectra (range 360–670 nm, resolution 0.38 nm) of IM Nor were obtained on Apr. 29.40 UT with the Blanco 4-m telescope (+ RC spectrograph) at Cerro Tololo. The spectrum is dominated by features characteristic of the nebular phase of development of a nova shell. Evident features include the Balmer series, He I 587.3- and 667.6-nm, He II 468.1nm, [N II] 575.3-nm, N III 463.7-nm, [O III] 496.1- and 500.6-nm, and [O I] 636.4-nm, as well as weaker emission lines due to [Fe II] and coronal lines such as [Fe III] 608.6-nm. Strong P-Cyg profiles (IAUC 7818) are not detected; however, many of the forbidden lines exhibit ‘castellate’ line profiles (Austin et al. 1996, A.J. 111, 869), suggesting the presence of multiple velocity components in the shell. The observed Hα line flux was ≈ 5.1 × 10^{−11} erg cm^{−2} s^{−1}. Follow-up observations with Chandra are anticipated, and additional ground-based observations of IM Nor are encouraged.

2002 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green